OpenStudy (anonymous):

how would you simplify this? 27=45(mod 8) so that its' residue is lesser than it's divisor (mod 8)

OpenStudy (kinggeorge):

I'm confused... Maybe I'm just not understanding what a residue is, but \[27 \neq 45 \quad (\text{\mod} \:\:\:8)\]Rather, \(27 = 3 \quad (\text{mod} \:\:\: 8) \) and \(45 = 5 \quad (\text{mod} \:\:\: 8) \)

OpenStudy (anonymous):

tha residue is 45, the other name for the residue is remainder. . i've been looking on the internet since then, but i haven't found any way on how to simplify such.

OpenStudy (anonymous):

excuse me, it's -27

OpenStudy (kinggeorge):

That makes more sense.

OpenStudy (anonymous):

can you help me?

OpenStudy (kinggeorge):

Well, the easiest way to simplify this problem, is to simply take\[{45 \over 8} = 5.625\]so you know the quotient is 5. Then\[45- 8*5 = 45-40 = 5\]so the residue/remainder would be 5. A similar process can be used for -27. In this case, both have residue of 5.

OpenStudy (anonymous):

you mean. . -27-40=45-40(mod8) would result to -67=5(mod 8)?

OpenStudy (kinggeorge):

That is correct.

OpenStudy (anonymous):

AHH, THANKS. .

OpenStudy (kinggeorge):

your welcome. Any other quick modular arithmetic questions?

OpenStudy (anonymous):

nope, at this moment. . perhaps, in other times. . lol thanks again. .